OpenStudy (anonymous):

How to write an equivalent expression using exponents for this question: sq. root, of the sq. root of 81x^8?

OpenStudy (anonymous):

\[\sqrt{}\sqrt{81x^{8}}\]

OpenStudy (anonymous):

and also, \[\sqrt{}\sqrt[3]{8x ^{7}}\]

OpenStudy (phi):

Do you know the answer if there was only 1 square root?

OpenStudy (anonymous):

no, I don't.. I get so confused with the square roots..

OpenStudy (phi):

\[\sqrt{stuff}= stuff^{\frac{1}{2}}\] try doing \[\sqrt{81 x^{8}}= ??^{\frac{1}{2}}\]

OpenStudy (anonymous):

ohh, it's 9^1/2

OpenStudy (phi):

half way... you also need (x^8)^(1/2)

OpenStudy (anonymous):

it says 4.

OpenStudy (phi):

And if you have something like\[(x^{8})^{\frac{1}{2}}= x^{8\cdot\frac{1}{2}}= x^{4}\] multiply the exponents

OpenStudy (anonymous):

ohhh... i see. So, after I half the base I do that?

OpenStudy (anonymous):

when there's two square roots?

OpenStudy (phi):

the answer with one square root is \[81^{\frac{1}{2}}x^{8\cdot\frac{1}{2}}= 9x^{4}\] so put the 9x^4 in a square root and do it again

OpenStudy (phi):

I'm not sure what they want for the final answer. If you rewrite the expression using exponents (and do not simplify it) it would look like this \[81^{\frac{1}{4}}x^{2}\]

OpenStudy (phi):

Your other problem is the square root of a cube root? a cube root is the same as using an exponent of 1/3 (instead of 1/2)

OpenStudy (anonymous):

ohhhhh, i see. Thanks!!